A force parallel to the x-axis applied to an object varies according to the linear graph (a straight line 45° of the x axis). Calculate the work done by the force when the object moves from x = 0 to x = 10. Find the equation of the line (F vs x) where F is the y-axis.


1
Expert's answer
2021-04-15T07:17:15-0400

"F(x)=x;"

"A=\int_0^{10}F(x)dx=\int _0^{10}xdx=\frac{x^2}{2}|_0^{10}=50~J."